simplify each expression by rewriting it as a single logarithm
log X+ log Y+ log Z=
4 ln M + 21n N=
6 log A- 3 log B=
X log A + Y log B- (Z log C + W log D)=
1
Expert's answer
2011-05-17T08:16:54-0400
log X + log Y + log Z = log XYZ
4 ln M + 2 ln N = ln M4N2
6 logA - 3 log B = log A6/B3
X log A + Y log B- (Z log C + W log D)= log AX BY - log CZ WD = log (AX BY)/(CZWD)